### v1

@@ 4 Revisit on The Fidelity Term of (5)
In order to develop an interactive segmentation model, it is necessary to modify the fidelity term in the segmentation model to incorporate user interaction. In this section, we thoroughly analyze the fidelity term. This analysis will facilitate our development of the interactive segmentation model.
@@
$E(c_{1},c_{2},\varphi):=\int_{\Omega}(I-J(c_{1},c_{2},\varphi))^{2}\mathrm{d}\bm{x}=\int_{\Omega}(I-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}.$ (7)
@@
$E(c_{1},c_{2},G)=\int_{\Omega}(I-c_{1}\mathcal{X}_{G}-c_{2}\mathcal{X}_{G^{c}})^{2}\mathrm{d}\bm{x}.$ (8)
A segmentation mask $G$ can divide an image domain into a foreground and a background. However, since the segmentation may be imperfect, there could be segmented regions that are not part of the true region of interest, which we call false positive. Those regions that are classified to be the background but are the actual region of interest are called false negative. To simplify our analysis, we can assume two cases: (1) the case with only false positive and (2) the case with only false negative. Thus, we can assume the segmentation result to have three domains, true positive (the correctly segmented foreground), true negative (the correctly segmented background), and false postive/false negative.
Figure 2: Notations for a three-colored image with three domains: $\Omega_{0}$ (background), $\Omega_{1}$ (foreground - wanted), and $\Omega_{2}$ (foreground - unwanted).
Following these considerations, as Figure 2, we set the image $I$ as follows
$I=\begin{cases}p_{0},&\bm{x}\in\Omega_{0},\\ p_{1},&\bm{x}\in\Omega_{1},\\ p_{2},&\bm{x}\in\Omega_{2}.\\ \end{cases}$ (9)
Since the domain $\Omega_{0}$ is the background region, minimizing the segmentation energy (8) can lead to three cases: ($E(c_{1},c_{2},G=\Omega_{1})$; ($E(c_{1},c_{2},G=\Omega_{2})$; (3)$E(c_{1},c_{2},G=\Omega_{1}\cup\Omega_{1})$. Denote the area of some domain $\Omega$ as $||\Omega||$. We can have the following theorem:
@@
The segmentation energy (8) for a three-colored image $I$ has only three local minimizes. They are
@@
This means that no partial of $\Omega_{i}$ can be included.
Proof 4.2.
To prove this, we need to check if there is a case that a partial of $\Omega_{0}$, $\Omega_{1}$ or $\Omega_{2}$ is included in a local minimum. To do so, we investigate the value of $E(G=\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2})$, where $\mathcal{D}_{0}\subset\Omega_{0}$, $\mathcal{B}_{0}=\Omega_{0}\setminus\mathcal{D}_{0}$, $\mathcal{D}_{1}\subset\Omega_{1}$, $\mathcal{B}_{1}=\Omega_{1}\setminus\mathcal{D}_{1}$, $\mathcal{D}_{2}\subset\Omega_{2}$ and $\mathcal{B}_{2}=\Omega_{2}\setminus\mathcal{D}_{2}$.
When $G=\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2}$, we have
$c_{1}=\frac{\int_{\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2}}I\mathrm{d}\bm{x}}{\int_{\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2}}1\mathrm{d}\bm{x}}=\frac{p_{0}||\mathcal{D}_{0}||+p_{1}||\mathcal{D}_{1}||+p_{2}||\mathcal{D}_{2}||}{||\mathcal{D}_{0}||+||\mathcal{D}_{1}||+||\mathcal{D}_{2}||}$ (10)
and
$c_{2}=\frac{\int_{\mathcal{B}_{0}\cup\mathcal{B}_{1}\cup\mathcal{B}_{2}}I\mathrm{d}\bm{x}}{\int_{\mathcal{B}_{0}\cup\mathcal{B}_{1}\cup\mathcal{B}_{2}}1\mathrm{d}\bm{x}}=\frac{p_{0}||\mathcal{B}_{0}||+p_{1}||\mathcal{B}_{1}||+p_{2}||\mathcal{B}_{2}||}{||\Omega_{0}||+||\mathcal{B}_{1}||+||\mathcal{B}_{2}||}.$ (11)
So we get
$E(c_{1},c_{2},G=\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2})$ $=\int_{\Omega}(I-c_{1}\mathcal{X}_{\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2}}-c_{2}\mathcal{X}_{\mathcal{B}_{0}\cup\mathcal{B}_{1}\cup\mathcal{B}_{2}})^{2}\mathrm{d}\bm{x}$ (12)
$=\int_{\Omega}((p_{0}-c_{1})\mathcal{X}_{\mathcal{D}_{0}}+(p_{1}-c_{1})\mathcal{X}_{\mathcal{D}_{1}}+(p_{2}-c_{1})\mathcal{X}_{\mathcal{D}_{2}}$
$\quad+(p_{0}-c_{2})\mathcal{X}_{\mathcal{B}_{0}}+(p_{1}-c_{2})\mathcal{X}_{\mathcal{B}_{1}}+(p_{2}-c_{2})\mathcal{X}_{\mathcal{B}_{2}})^{2}\mathrm{d}\bm{x}$
$=(p_{0}-c_{1})^{2}||\mathcal{D}_{0}||+(p_{1}-c_{1})^{2}||\mathcal{D}_{1}||+(p_{2}-c_{1})^{2}||\mathcal{D}_{2}||$
$E(c_{1},c_{2},G)$ $=\frac{((p_{0}-p_{1})D_{1}+(p_{0}-p_{2})D_{2})^{2}}{(D_{0}+D_{1}+D_{2})^{2}}D_{0}+\frac{((p_{1}-p_{0})D_{0}+(p_{1}-p_{2})D_{2})^{2}}{(D_{0}+D_{1}+D_{2})^{2}}D_{1}$ (13)
$\quad+\frac{((p_{2}-p_{0})D_{0}+(p_{2}-p_{1})D_{1})^{2}}{(D_{0}+D_{1}+D_{2})^{2}}D_{2}+\frac{((p_{0}-p_{1})B_{1}+(p_{0}-p_{2})B_{2})^{2}}{(B_{0}+B_{1}+B_{2})^{2}}B_{0}$
$\quad+\frac{((p_{1}-p_{0})B_{0}+(p_{1}-p_{2})B_{2})^{2}}{(B_{0}+B_{1}+B_{2})^{2}}B_{1}+\frac{((p_{2}-p_{0})B_{0}+(p_{2}-p_{1})B_{1})^{2}}{(B_{0}+B_{1}+B_{2})^{2}}B_{2}$
$=\frac{d_{12}^{2}D_{1}D_{2}+d_{01}^{2}D_{0}D_{1}+d_{02}^{2}D_{0}D_{2}}{D_{0}+D_{1}+D_{2}}+\frac{d_{12}^{2}B_{1}B_{2}+d_{01}^{2}B_{0}B_{1}+d_{02}^{2}B_{0}B_{2}}{B_{0}+B_{1}+B_{2}}$
$=d_{12}^{2}(D_{2}-\frac{D_{0}D_{2}+D_{2}^{2}}{D_{0}+D_{1}+D_{2}})+d_{01}^{2}(D_{0}-\frac{D_{0}^{2}+D_{0}D_{2}}{D_{0}+D_{1}+D_{2}})+d_{02}^{2}\frac{D_{0}D_{2}}{D_{0}+D_{1}+D_{2}}$
$\quad+d_{12}^{2}(B_{2}-\frac{B_{0}B_{2}+B_{2}^{2}}{B_{0}+B_{1}+B_{2}})+d_{01}^{2}(B_{0}-\frac{B_{0}^{2}+B_{0}B_{2}}{B_{0}+B_{1}+B_{2}})+d_{02}^{2}\frac{B_{0}B_{2}}{B_{0}+B_{1}+B_{2}}$
$\quad+d_{02}^{2}\frac{D_{0}D_{2}}{D_{0}+A_{1}+D_{2}-B_{1}}$
Treating $B_{0}$, $B_{1}$, $B_{2}$ as three variables, we can have $B_{0}\in[0,A_{0}]$, $B_{1}\in[0,A_{1}]$, $B_{2}\in[0,A_{2}]$ as $\mathcal{B}_{0}\subset\Omega_{0}$, $\mathcal{B}_{1}\subset\Omega_{1}$, $\mathcal{B}_{2}\subset\Omega_{2}$. Naturally, $D_{0}=A_{0}-B_{0}$, $D_{1}=A_{1}-B_{1}$, $D_{2}=A_{2}-B_{2}$. Further, define a function:
$F(B_{0},B_{1},B_{2})=E(c_{1},c_{2},G=\mathcal{D}_{0}\cup\mathcal{D}_{1}\cup\mathcal{D}_{2}).$ (14)
We then compute $\frac{\partial F}{\partial B_{1}}$:
$\frac{\partial F}{\partial B_{1}}$ $=-\frac{d_{12}^{2}(D_{0}D_{2}+D_{2}^{2})+d_{01}^{2}(D_{0}^{2}+D_{0}D_{2})-d_{02}^{2}D_{0}D_{2}}{(D_{0}+A_{1}+D_{2}-B_{1})^{2}}$ (15)
$\begin{split}&d_{12}^{2}(B_{0}B_{2}+B_{2}^{2})+d_{01}^{2}(B_{0}^{2}+B_{0}B_{2})-d_{02}^{2}B_{0}B_{2}\\ \geq&d_{12}^{2}B_{0}B_{2}+d_{01}^{2}B_{0}B_{2}+2d_{12}d_{01}B_{0}B_{2}-d_{02}^{2}B_{0}B_{2}\\ =&(d_{01}+d_{12})^{2}B_{0}B_{2}-d_{02}^{2}B_{0}B_{2}\geq 0\end{split}$
and
$\begin{split}&d_{12}^{2}(D_{0}D_{2}+D_{2}^{2})+d_{01}^{2}(D_{0}^{2}+D_{0}D_{2})-d_{02}^{2}D_{0}D_{2}\\ \geq&d_{12}^{2}D_{0}D_{2}+d_{01}^{2}D_{0}D_{2}+2d_{12}d_{01}D_{0}D_{2}-d_{02}^{2}D_{0}D_{2}\\ =&(d_{01}+d_{12})^{2}D_{0}D_{2}-d_{02}^{2}D_{0}D_{2}\geq 0.\end{split}$
As $\frac{\partial F}{\partial B_{1}}$ is monotonously decreasing for any given $B_{0}$ and $B_{2}$, there is at most one maximum within $B_{1}\in(0,A_{1})$ for any given $B_{0}$ and $B_{2}$. The minimum of $F(B_{0},B_{1},B_{2})$ must only be obtained when $B_{1}=0$ or $B_{1}=A_{1}$ for any given $B_{0}$ and $B_{2}$. So is for $B_{2}$ given any $B_{0}$ and $B_{1}$. Thus, the minimum can never be obtained when partial of the domain $\Omega_{i}$ is included.
@@
Next, we compute the corresponding energy value with respect to these different cases. When $G=\Omega_{1}$, we have $c_{1}=\frac{\int_{\Omega_{1}}I\mathrm{d}\bm{x}}{\int_{\Omega_{1}}1\mathrm{d}\bm{x}}=p_{1}$ and $c_{2}=\frac{\int_{\Omega_{1}^{c}}I\mathrm{d}\bm{x}}{\int_{\Omega_{1}^{c}}1\mathrm{d}\bm{x}}=\frac{p_{2}||\Omega_{2}||+p_{0}||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||}$. Then, we can get
$E(c_{1},c_{2},G=\Omega_{1})$ $=\int_{\Omega}(I-c_{1}\mathcal{X}_{\Omega_{1}}-c_{2}\mathcal{X}_{\Omega_{2}\cup\Omega_{0}})^{2}\mathrm{d}\bm{x}$ (16)
$=\int_{\Omega}(p_{0}\mathcal{X}_{\Omega_{0}}+p_{1}\mathcal{X}_{\Omega_{1}}+p_{2}\mathcal{X}_{\Omega_{2}}-p_{1}\mathcal{X}_{\Omega_{1}}-c_{2}\mathcal{X}_{\Omega_{0}}-c_{2}\mathcal{X}_{\Omega_{2}})^{2}\mathrm{d}\bm{x}$
$=\int_{\Omega}|(p_{0}-c_{2})\mathcal{X}_{\Omega_{0}}+(p_{2}-c_{2})\mathcal{X}_{\Omega_{2}}|^{2}\mathrm{d}\bm{x}$
$=(p_{0}-c_{2})^{2}\int_{\Omega}|\mathcal{X}_{\Omega_{0}}|^{2}\mathrm{d}\bm{x}+(p_{2}-c_{2})^{2}\int_{\Omega}|\mathcal{X}_{\Omega_{2}}|^{2}\mathrm{d}\bm{x}$
@@
$=\left(\frac{p_{0}||\Omega_{2}||-p_{2}||\Omega_{2}||}{||\Omega_{2}||+||\Omega_{0}||}\right)^{2}||\Omega_{0}||+\left(\frac{p_{2}||\Omega_{0}||-p_{0}||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||}\right)^{2}||\Omega_{2}||$
@@
Similarly, when $G=\Omega_{2}$, we have $c_{1}=\frac{\int_{\Omega_{2}}I\mathrm{d}\bm{x}}{\int_{\Omega_{2}}1\mathrm{d}\bm{x}}=p_{2}$, $c_{2}=\frac{\int_{\Omega_{2}^{c}}I\mathrm{d}\bm{x}}{\int_{\Omega_{2}^{c}}1\mathrm{d}\bm{x}}=\frac{p_{1}||\Omega_{1}||+p_{0}||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}$, and
$E(c_{1},c_{2},G=\Omega_{2})$ $=\int_{\Omega}(I-c_{1}\mathcal{X}_{\Omega_{2}}-c_{2}\mathcal{X}_{\Omega_{1}\cup\Omega_{0}})^{2}\mathrm{d}\bm{x}$ (17)
$=(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}.$
And for $G=\Omega_{1}\cup\Omega_{2}$, we get $c_{1}=\frac{\int_{\Omega_{1}\cup\Omega_{2}}I\mathrm{d}\bm{x}}{\int_{\Omega_{1}\cup\Omega_{2}}1\mathrm{d}\bm{x}}=\frac{p_{1}||\Omega_{1}||+p_{2}||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}$, $c_{2}=\frac{\int_{\Omega_{0}}I\mathrm{d}\bm{x}}{\int_{\Omega_{0}}1\mathrm{d}\bm{x}}=p_{0}$, and
$E(c_{1},c_{2},G=\Omega_{1}\cup\Omega_{2})$ $=\int_{\Omega}(I-c_{1}\mathcal{X}_{\Omega_{1}\cup\Omega_{2}}-c_{2}\mathcal{X}_{\Omega_{0}})^{2}\mathrm{d}\bm{x}$ (18)
$=(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}.$
For a clear three-colored image, we can optimize the energy functional (8) and obtain $G^{*}$ by comparing the energy values of the above three cases:
@@
    $\begin{cases}(p_{0}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{2}||+||\Omega_{0}||}\leq(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{0}||},\\ (p_{0}-p_{2})^{2}\frac{||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||}\leq(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{2}||}.\end{cases}$ (19)
@@
    $\begin{cases}(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{0}||}\leq(p_{0}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{2}||+||\Omega_{0}||},\\ (p_{0}-p_{1})^{2}\frac{||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}\leq(p_{1}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}.\end{cases}$ (20)
@@
    $\begin{cases}(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{2}||}\leq(p_{0}-p_{2})^{2}\frac{||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||},\\ (p_{1}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}\leq(p_{0}-p_{1})^{2}\frac{||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}.\end{cases}$ (21)
@@
The minimizer $G^{*}$ for the segmentation energy (8) for a clear three-colored image $I$ is one of the following three:
@@
    $G^{*}=\Omega_{1}$, when (19) is satisfied;
@@
    $G^{*}=\Omega_{2}$, when (20) is satisfied;

### v2

@@ 4.3.1 Revisit on the fidelity term of (3)
In this part, we thoroughly analyze the fidelity term of (3). This analysis will facilitate our development of the interactive segmentation model.
@@
(6) $E(c_{1},c_{2},\varphi):=\int_{\Omega}(I-J(c_{1},c_{2},\varphi))^{2}\mathrm{d}\bm{x}=\int_{\Omega}(I-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}.$
@@
(7) $E(c_{1},c_{2},G)=\int_{\Omega}(I-c_{1}\mathcal{X}_{G}-c_{2}\mathcal{X}_{G^{c}})^{2}\mathrm{d}\bm{x}.$
A segmentation mask $G$ can divide an image domain into a foreground and a background. However, since the segmentation may be imperfect, there could be segmented regions that are not part of the true region of interest, which we call false positive. Those regions that are classified to be the background but are the actual region of interest are called false negative. To simplify our analysis, we can assume two cases: (1) the case with only false positive and (2) the case with only false negative. Thus, we can assume the segmentation result to have three domains, true positive (the correctly segmented foreground), true negative (the correctly segmented background), and false positive/false negative.
Figure 4: Notations for a three-value image with three domains: $\Omega_{0}$ (background), $\Omega_{1}$ (foreground - wanted), and $\Omega_{2}$ (foreground - unwanted).
Following these considerations, as Figure 4, we set a three-value image $I$ as follows
(8) $I=\begin{cases}p_{0},&\bm{x}\in\Omega_{0},\\ p_{1},&\bm{x}\in\Omega_{1},\\ p_{2},&\bm{x}\in\Omega_{2},\\ \end{cases}$
where the domain $\Omega_{0}$ is the background region. Denote the area of some domain $\Omega$ as $||\Omega||$. Then the following theorem indicates that minimizing the segmentation energy (7) can only lead to three cases: ($E(c_{1},c_{2},G=\Omega_{1})$; ($E(c_{1},c_{2},G=\Omega_{2})$; (3)$E(c_{1},c_{2},G=\Omega_{1}\cup\Omega_{1})$.
@@
The segmentation energy (7) for a three-value image $I$ only has three local minimizes. They are
@@
This means that no subset of $\Omega_{i}$ can be included.
Proof 4.3.
See Appendix A.
@@
Next, we compute the corresponding energy value with respect to these different cases. When $G=\Omega_{1}$, we have $c_{1}=p_{1}$ and $c_{2}=\frac{p_{2}||\Omega_{2}||+p_{0}||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||}$. Then, we can get
(9) $E(c_{1},c_{2},G=\Omega_{1})$ $=\int_{\Omega}(I-c_{1}\mathcal{X}_{\Omega_{1}}-c_{2}\mathcal{X}_{\Omega_{2}\cup\Omega_{0}})^{2}\mathrm{d}\bm{x}$
@@
Similarly, when $G=\Omega_{2}$, we have $c_{1}=p_{2}$, $c_{2}=\frac{p_{1}||\Omega_{1}||+p_{0}||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}$, and
(10) $E(c_{1},c_{2},G=\Omega_{2})=(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}.$
And for $G=\Omega_{1}\cup\Omega_{2}$, we get $c_{1}=\frac{p_{1}||\Omega_{1}||+p_{2}||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}$, $c_{2}=p_{0}$, and
(11) $E(c_{1},c_{2},G=\Omega_{1}\cup\Omega_{2})=(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}.$
Hence, for a clear three-value image, we can optimize the energy functional (7) and obtain $G^{*}$ by comparing the energy values of the above three cases:
@@
    (12) $\begin{cases}(p_{0}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{2}||+||\Omega_{0}||}\leq(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{0}||},\\ (p_{0}-p_{2})^{2}\frac{||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||}\leq(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{2}||}.\end{cases}$
@@
    (13) $\begin{cases}(p_{0}-p_{1})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{0}||}\leq(p_{0}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{2}||+||\Omega_{0}||},\\ (p_{0}-p_{1})^{2}\frac{||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}\leq(p_{1}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}.\end{cases}$
@@
    (14) $\begin{cases}(p_{1}-p_{2})^{2}\frac{||\Omega_{1}||}{||\Omega_{1}||+||\Omega_{2}||}\leq(p_{0}-p_{2})^{2}\frac{||\Omega_{0}||}{||\Omega_{2}||+||\Omega_{0}||},\\ (p_{1}-p_{2})^{2}\frac{||\Omega_{2}||}{||\Omega_{1}||+||\Omega_{2}||}\leq(p_{0}-p_{1})^{2}\frac{||\Omega_{0}||}{||\Omega_{1}||+||\Omega_{0}||}.\end{cases}$
@@
The minimizer $G^{*}$ of the segmentation energy (7) for a clear three-value image $I$ is one of the following three:
@@
    when (12) is satisfied, $G^{*}=\Omega_{1}$;
@@
    when (13) is satisfied, $G^{*}=\Omega_{2}$;